OpenStudy (anonymous):

Find all the values of r such that y=Ax^r satisfies the equation: y''+(x^-1)y'-y^3=0 Give a reason why other solutions must exist for this equation.

OpenStudy (anonymous):

Hello estudier

OpenStudy (anonymous):

That's an improvement on the previous version, lol

OpenStudy (anonymous):

Hi..

OpenStudy (anonymous):

I can get to \[r(r-1)Ax^{r-2} + rAx^{r-2}-A^{3}x{3r}=0\]

OpenStudy (anonymous):

By differentiating y=Ax^r and plugging in

OpenStudy (anonymous):

Should be this actually \[r(r-1)Ax^{r-2} + rAx^{r-2}-A^{3}x^{3r}=0\]

OpenStudy (anonymous):

No one helped u...post it again.

OpenStudy (anonymous):

Just that last bit, solve for r....